OpenStudy (ashtarothfick):

HELP! A drug trial is testing the effectiveness of two drugs. If 50 patients are given Drug A, 20 patients are given Drug B, and 100 patients are given a placebo, what is the probability that a patient will NOT be given a placebo? A. 5/17 B. 10/17 C. 2/17 D. 7/17

OpenStudy (kropot72):

The total number of patients in the trial is 170. The number of patients who are not given a placebo is 50 + 20 = 70. Therefore the required probability is 70/170 = 7/17.

OpenStudy (ashtarothfick):

Thank you so much I appreciate. you you possibly help me with five others?

OpenStudy (kropot72):

You're welcome :) Please post the others as separate questions. Thanks.